A perimeter of a triangle is 49 cm. The first side is 2 cm less than twice the second side,and the third side is 3 cm longer than the second side. Find the lenght of each side.
Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39838)

x, y, z the sides.
perimeter, 49=x+y+z.
x=y-2, z=y+3.

Substitute into perimeter equation.
(y-2)+y+(y+3)=49
Solve for y and compute the other sides x and z.

Answer by MathTherapy(10858)


A perimeter of a triangle is 49 cm. The first side is 2 cm less than twice the second side,and the third side is 3 cm longer than the second side. Find the lenght of each side.


Let length of second side be S
Then length of first side is: 2S - 2, and length of third side is: S + 3
Thus, S + 2S - 2 + S + 3 = 49
Solve this to determine S, or the second side. Thereafter, make the necessary substitutions to find the other two sides.
Then do a check!!
Send comments, “thank-yous,” and inquiries to “D” at MathMadEzy@aol.com.
Further help is available, online or in-person,for a fee, obviously.
For FREE info and answers to questions about the ASVAB exam, the NYS 3 – 8 city/state wide exams,
GENERAL MATH QUESTIONS, MATH HOMEWORK QUESTIONS, or MATH QUESTIONS related to the Regents Integrated Algebra,
Regents Geometry, Regents Algebra 2/Trigonometry, SHSAT, COOP/HSPT/TACHS, PSAT, SAT, ACT, SSAT/ISEE, GRE, CLEP,
and the GED, you can visit: http://asvabstudyzone.freeforums.net/.